OpenStudy (anonymous):

Which expression is a cube root of -1+i sqrt3

OpenStudy (anonymous):

(cos(40) + isin(40)) (cos(120) + isin(120)) (cos(280) + isin(280)) (cos(320) + isin(320))

OpenStudy (anonymous):

\[\sqrt{x^2+y^2} e ^{\tan^{-1} y/x}\]

OpenStudy (kropot72):

None of the expressions given is the solution, the reason being that when \[-1+i \sqrt{3}\] is converted to polar form r = 2. But when the cube root of the complex number is taken r will become \[\sqrt[3]{2}\] which is not in any of the choices.

OpenStudy (anonymous):

z=-1+i sqrt(3) ||z||=sqrt(1+3)=2 z=2 (-0.5+i sqrt(3)/2) cos(theta)=-0.5 .............. sin(theta)=sqrt(3)/2 -------> theta=120 z=2 exp(120 i ) \[\sqrt[3]{z}=\sqrt[3]{2}*e ^{(120/3+2k pi)i}\] k=0,1,2 then \[\sqrt[3]{z}=\sqrt[3]{2}*e ^{40i}=\sqrt[3]{2}*(\cos 40+ i \sin 40)\]